OpenStudy (kkutie7):

two trains start 5 min apart. Starting from rest, each is capable of a maximum speed of 300 km/h after uniformly accelerating over a distance of 4 km. (a) What is the acceleration of each train? (b) how far is the first train ahead when the second starts? (c) How far apart are they when they are both traveling at maximum speed?

OpenStudy (kkutie7):

I can only use these equations: \[v=v_{0}+at\] \[r=r_{0}+v_{0}t+\frac{1}{2}at^{2}\]

OpenStudy (anonymous):

i really should be quiet because i have no idea how to do this, but after reading it 3 times it strikes me that something is missing how in the world are you supposed to know the acceleration?

OpenStudy (kkutie7):

I'm not sure haha. I'm trying to rack my brain for something...

OpenStudy (anonymous):

are you sure there is not some missing part ? like "the meet one hour later" or something?

OpenStudy (anonymous):

i have two cars they do not accelerate at the same rate if i ask you for the rate, you cannot tell me

OpenStudy (kkutie7):

No I wrote it word for word from my work and it tends to be right so I'm not sure.

OpenStudy (anonymous):

ooh maybe this line tarting from rest, each is capable of a maximum speed of 300 km/h after uniformly accelerating over a distance of 4 km. means that it takes 4 km to accelerate to a speed of 300

OpenStudy (anonymous):

\[v=v_{0}+at\] if the start at rest then it is \[v=at\]

OpenStudy (kkutie7):

yeah I'm trying to work with that but I keeps messing up I have the answer: \[0.87 m/s^{2}\]

OpenStudy (anonymous):

i have an idea (could be wrong)

OpenStudy (anonymous):

\(v_0=0\) since they start at rest and also \(r_0=0\)

OpenStudy (anonymous):

then \[300=at\]and \[\frac{1}{2}at^2=4\]

OpenStudy (anonymous):

we can solve that for \(t\) then for \(a\)

OpenStudy (anonymous):

\[\frac{.5at^2}{at}=\frac{4}{300}\implies t=\frac{8}{300}\]

OpenStudy (anonymous):

no that doesn't seem to work, damn

OpenStudy (kkutie7):

maybe start by doing this? \[4km*\frac{1000m}{1km}=4000m\] \[\frac{300km}{h}*\frac{1000m}{1km}=\frac{300000m}{h}\]

OpenStudy (anonymous):

everything is given in terms of km i would keep those units until the end but i did not notice that your answer was in terms of meter, so lets see if what i wrote is correct

OpenStudy (anonymous):

nope it doesn't even make sense

OpenStudy (kkutie7):

aww man lol

OpenStudy (anonymous):

isn't the internet wonderful? i wish they had this when i was in school https://answers.yahoo.com/question/index?qid=20110125211241AAqffOy

OpenStudy (kkutie7):

is there a way to solve for one equation and put it into the other

OpenStudy (kkutie7):

awesome thank you!

OpenStudy (anonymous):

yw